# Better Aligned with Survey Respondents or Training Data? Unveiling Political Leanings of LLMs on U.S. Supreme Court Cases

### v2

@@ Abstract
The increased adoption of Large Language Models (LLMs) and their potential to shape public opinion have sparked interest in assessing these models’ political leanings. Building on previous research that compared LLMs and human opinions and observed political bias in system responses, we take a step further to investigate the underlying causes of such biases by empirically examining how the values and biases embedded in training corpora shape model outputs. Specifically, we propose a method to quantitatively evaluate political leanings embedded in the large pretraining corpora. Subsequently we investigate to whom are the LLMs’ political leanings more aligned with, their pretrainig corpora or the surveyed human opinions. As a case study, we focus on probing the political leanings of LLMs in 32 U.S. Supreme Court cases, addressing contentious topics such as abortion and voting rights. Our findings reveal that LLMs strongly reflect the political leanings in their training data, and no strong correlation is observed with their alignment to human opinions as expressed in surveys. These results underscore the importance of responsible curation of training data and the need for robust evaluation metrics to ensure LLMs’ alignment with human-centered values.
@@
Shanshan Xu1, Santosh T.Y.S.S1,Yanai Elazar2,3 Quirin Vogel1, Barbara Plank4,5, Matthias Grabmair1 1Technical University of Munich, Germany 2Allen Institute for AI 3University of Washington 4IT University of Copenhagen, Denmark 5LMU Munich & Munich Center for Machine Learning (MCML)
@@ 1 Introduction
Values and opinions expressed by large language models (LLMs) such as ChatGPT OpenAI (2023) have the potential to influence public perceptions and shape societal views. As such, it is essential to evaluate the political leanings of LLMs. Previous research (e.g., Perez et al. 2023; Hartmann et al. 2023) found a misalignment between LLM outputs and human opinions, and several identified a left-leaning bias of LLMs.
@@
LLMs derive their knowledge primarily from their pre-training data, which are typically composed of internet text. However, the structure of online content regularly skews toward overrepresenting specific viewpoints Galeazzi et al. (2024), often introducing biases into these models. Such biases can then propagate through model outputs, reinforcing existing societal biases Feng et al. (2023). Despite increasing awareness of these issues, to the best of our knowledge, no prior work has shown how to quantitatively analyze political bias in the content of large pre-training corpora and subsequently assess how well LLM outputs align with human opinions versus the biases in training data. To address this gap, we propose a pipeline to retrieve relevant documents from the pretraining corpora, then evaluate the political leanings expressed in these documents, and subsequently assess the alignment of political leanings in pretraining corpora with the responses generated by the LLM.
As a case study, we focus on US Supreme Court cases, which frequently address contentious and politically charged issues, such as death penalty, abortion, same-sex marriage, and voting rights, making them strong indicators of political leanings. Leveraging the SCOPE Jessee et al. (2022) [^1] Jessee et al. 2022 has not named the dataset. Hereafter, we refer to the dataset as SCOPE: Supreme COurt Case Political Evaluation. survey data on US Supreme Court cases from political studies, this paper examines the political leanings of eight LLMs and five open-source pre-training corpora, comparing them to human survey responses and Supreme Court rulings [^2] We will release the code and data upon publication.. The main contributions of our work are threefold:
@@
  • Our empirical findings indicate that LLMs exhibit significant alignment with their training corpora, yet we do not find strong alignment with human opinions. This highlights the critical need for responsible data curation and rigorous evaluation of LLM alignment.
@@ 2 Background
Evaluating LLM-Human Alignment
Recent research has increasingly focused on probing LLMs political opinions. Most approaches typically follow a two-stage process: (1) assessing an LLM’s political stance on specific topics, and (2) measuring how closely its responses align with human opinions. A common strategy for evaluating LLM opinions involves using political orientation tests (e.g., Political Compass Test333www.politicalcompass.org/test, as in Röttger et al. 2024; Feng et al. 2023) or survey questionnaires (e.g., PewResearch ATP444www.pewresearch.org/writing-survey-questions/, as in Santurkar et al. 2023). To quantify the alignment between human and LLM responses, prior work typically measures the similarity of their opinion distributions using either (a) distance-based metrics—such as the 1-Wasserstein distance Santurkar et al. (2023); Sanders et al. (2023) and Jensen–Shannon divergence Durmus et al. (2024)—or (b) statistical analyses, including Cohen’s Kappa Argyle et al. (2023); Hwang et al. (2023) and Pearson correlation coefficients Movva et al. (2024). We refer to Ma et al. 2024 for an extensive survey of methods in this area. In this work, we use SCOPE to probe LLMs’ political opinions because it offers several advantages over the above-mentioned political surveys used in previous studies: The cases in SCOPE are selected by experts, ensuring that they address the most important and publicly salient legal topics. Experts carefully word the questions and response options to be understandable to the general public. Moreover, political experts have annotated each case with thoughtfully chosen keywords, which facilitate our retrieval of relevant documents from large pretraining corpora, as detailed in Sec 4.1.3.
LLMs and their pretraining corpra
@@
Existing studies have explored the impact of biases in training corpora on LLM behavior, primarily through second-stage controlled training setups such as continual pretraining Feng et al. (2023); Chalkidis and Brandl (2024). While continual pretraining can offer valuable insights into the causal links between training data and model outputs, these studies rarely applied to study LLMs’ behavior based on initial pretraining phase, where biases are fundamentally embedded. Additionally, it is also computationally expensive to conduct such extensive pre-training experiments on initial phase. An alternative strategy involves investigating the correlation between biases in training corpora and those in model outputs Seshadri et al. (2024). Previously this approach has been under exploited, primarily due to the limited accessibility of large-scale pretraining datasets. Many commercial LLM providers (e.g., GPT-4 by OpenAI 2023 and Claude by Anthropic 2023) disclose minimal information about their training sets, not even corpus size or data source. With the growing call in the academic community for transparency and accessibility of LLM pretraining data Pistilli et al. (2023); McMillan-Major et al. (2024), several organizations have begun to make large-scale pretraining datasets publicly available, including RedPajama (Weber et al., 2024) and Dolma Soldaini et al. (2024). These initiatives are complemented by the development of APIs and analytical tooling platforms, such as WIMDB Elazar et al. (2024), which facilitate comprehensive analysis of the corpora. In this paper, we leverage WIMDB to analyze the political leanings in five publicly accessible corpora and subsequently evaluate how these leanings correlate with the outputs generated by various LLMs.
@@ 3.1 Dataset
In political science, researchers often estimate individuals’ or groups’ political preferences and ideological positions by analyzing observable behaviors, such as voting patterns and survey responses Martin and Quinn (2002); Ho et al. (2008). For example, Jessee et al. (2022) created SCOPE to gather respondents’ views on Supreme Court decisions. By comparing collected survey responses with the Court’s voting record, they demonstrated that the Court has adopted a more conservative stance than the general U.S. public. In this study, we use SCOPE Jessee et al. (2022) to prompt various LLMs to assess their political leanings and subsequently compare their alignment with surveyed human opinions and political leanings in their training data.
The SCOPE dataset comprises of 32 cases, each represented by a binary-choice question asking respondents to express their views on the Court’s ruling as either supportive (pro) or opposing (opp). Fig 2 provides an example of a survey question. Tab 2 in App C lists all 32 cases along with their corresponding legal topics in the SCOPE dataset. For each case, between 1,500 and 2,158 respondents indicate whether they are pro or opp regarding the Court’s decision. Additionally, SCOPE captures each respondent’s self-identified political ideology, enabling the categorization of participants into self-identified Democrats or Republicans. Tab 2 in App C showcases the distribution of choices {pro, opp} among the overall surveyed respondents, as well as within the self-identified Democratic and Republican respondents. Further descriptive statistics on respondents’ backgrounds are available in the original study Jessee et al. (2022).
@@
Table 1: Details of evaluated LLMs, along with the dataset that the model trained on. * signifies that the model was not trained exactly on this dataset, due to filtering, using additional data, or the original data being private.
@@ 3.3 Pretraining Corpora
Tab 1 lists the corresponding pretraining corpora (when available) of the LLMs we investigated in this work. It is important to note that among the various LLM-corpora pairs we consider, only the OLMo-SFT and OLMo-Instruct models were trained directly on the pretraining corpus Dolma Soldaini et al. (2024). While for all other pairs, the LLMs may not have been trained exactly on the versions of the corpora we consider, due to factors such as filtering, or inclusion of additional data Elazar et al. (2024). Despite these discrepancies, we treat the documented corpora as reasonable proxies for analysis, as they represent the closest publicly available approximations of the actual training data for these models [^5] All models examined in this paper have undergone post-training, such as Supervised or Instruction Fine-Tuning, which may also influence the opinions in models outputs. However, prior research Feng et al. (2023) suggests that the shift introduced by post-training is relatively small. We also explored the correlation between LLMs’ political leanings and that in their post-training data, but did not observe any significant correlation. Further discussions can be found in App G..
@@ 4 Methodology
We employ a three-stage process to examine LLMs’ political leanings and further compare their alignment with surveyed human opinions and their pretraining corpora, when available. First, we introduce how we assess the political leanings of different entities [^6] We use entity to refer to either a group of surveyed respondents, Supreme Court justices, LLM-generated responses, or content within the training data.. Next, we measure the political leanings alignment among them in Sec 4.2. Finally, we conduct significance tests to determine whether the observed differences in LLM alignment with different entities are statistically significant in Sec 4.3.
@@ 4.1 Extracting the Preference Distributions
In this section, we outline the methodology used to extract the preference distribution of various entities. We divide the entities to three categories - the humans $D_{H}$, the LLMs $D_{M}$, and the pretraining corpora $D_{C}$.
4.1.1 The Humans
Under the category of the humans, we consider the preference distribution of four entities: $D_{H}=\{D_{\mathrm{pub}},D_{\mathrm{dem}},D_{\mathrm{rep}},D_{\mathrm{court}}\}$. Here, $D_{\mathrm{pub}}$ represents the preference distribution of the overall surveyed respondents, while $D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$ correspond to surveyed self-identified Democrat and Republican respondents, respectively; $D_{\mathrm{court}}$ represents the preference distribution of the Court. All preference distributions are Bernoulli, with the respective parameter estimated from the data. For the $D_{\mathrm{court}}$, we fetch the judges’ votes from the Supreme Court Database Spaeth et al. (2024) [^7] http://scdb.wustl.edu/ , and then calculate $D_{\mathrm{court}}$ as the ratio of justices who agree (pro) / dissent (opo) with the majority decision. For $D_{pub},D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$ we calculate them as the ratios of {pro} versus {opo} to the court’s decisions among the respondents based on data retrieved from SCOPE.
4.1.2 The LLMs
Under the category of the LLMs $D_{M}$, we probe the political preferences of eight LLMs as listed in Tab 1. Following Scherrer et al. 2023, for each survey case in SCOPE, we created six different prompt templates, and we then sample five responses from each of the six prompt variations from the LLMs at a temperature setting of 1, yielding in a total of 30 responses per case per model. Example prompt templates and detailed prompt creation process can be found in Fig 6 in App B.
To map the LLM-generated answers to one of the given choice options, we employed an iterative, rule-based matching pipeline, as explained in App B. The preference distribution, denoted as $D_{m}=p_{m}(a_{j}\mid q_{i})$, reflects the ratio of support versus opposition to the court’s decision across the 30 generated responses for model $m$ on case $q_{i}$, illustrated in Fig 6 in App B.
4.1.3 The Training Data
@@
For each of the 32 cases $q_{i}$ in the SCOPE survey, we compile a set of keywords $K_{i}$ to retrieve relevant documents $T_{i}$ from the pretraining corpora using the WIMDB API Elazar et al. (2024)- a tool designed to facilitate analysis of large-scale pretraining corpora. For example, in the case Baze v. Rees [^8] Baze v. Rees, 553 U.S. 35 (2008), addresses whether lethal injection for executions was constitutional or not., we use keywords such as [lethal injection; capital punishment; Baze; Rees], retrieving 206 documents from the Dolma corpus. Further details on keyword selection and retrieval statistics for each case are provided in App C. Additionally, an example of a retrieved document is included in App I.
@@
We use Llama3-70B to assess the political stance $s_{i}^{j}$ of each retrieved document $t_{i}^{j}\in T_{i}$. The model is prompted to evaluate the document’s level of support for the court’s decision on a Likert scale from 1 (strongly against) to 5 (strongly supportive). If a document is unrelated to the case’s political issue, the model is instructed to return ‘Not related.’
@@
Quality Assessment of Stance Detection To evaluate the reliability of Llama370B’s stance detection, we manually annotated a randomly selected sample of 80 retrieved documents. We measure the agreement between human and model labels using Spearman’s rank correlationSpearman (1904). The overall Spearman’s $\rho$ is 0.68, indicating good alignment between Llama370B and human annotators. App C offers details on the quality assessment process. To evaluate the robustness of our document retrieval method, we performed a bootstrapping analysis by iteratively excluding 20% of retrieved documents. This procedure revealed no significant shifts in measured political leanings (see App C for methodological details). Although differences in keyword selection may affect document retrieval and thereby influence corpus-level political stance estimates, our findings demonstrate that results are resilient to changes in the retrieved documents.

### v3

@@ Abstract
Recent works have shown that Large Language Models (LLMs) have a tendency to memorize patterns and biases present in their training data, raising important questions about how such memorized content influences model behavior. One such concern is the emergence of political bias in LLM outputs. In this paper, we investigate the extent to which LLMs’ political leanings reflect memorized patterns from their pretraining corpora. We propose a method to quantitatively evaluate political leanings embedded in the large pretraining corpora. Subsequently we investigate to whom are the LLMs’ political leanings more aligned with, their pretrainig corpora or the surveyed human opinions. As a case study, we focus on probing the political leanings of LLMs in 32 U.S. Supreme Court cases, addressing contentious topics such as abortion and voting rights. Our findings reveal that LLMs strongly reflect the political leanings in their training data, and no strong correlation is observed with their alignment to human opinions as expressed in surveys. These results underscore the importance of responsible curation of training data, and the methodology for auditing the memorization in LLMs to ensure human-AI alignment.
@@
Shanshan Xu1, Santosh T.Y.S.S1,Yanai Elazar2,3 Quirin Vogel1, Barbara Plank4, Matthias Grabmair1 1Technical University of Munich, Germany 2Allen Institute for AI 3University of Washington 4LMU Munich & Munich Center for Machine Learning (MCML)
@@ 1 Introduction
LLMs derive their knowledge primarily from their pre-training data, which are typically composed of internet text. These sources, however, tend to overrepresent certain perspectives and ideologies, leading to biased training distributions Galeazzi et al. (2024). Previous work reveals that LLMs tend memorize parts of their training data Carlini et al. (2021). As a result, LLMs risk memorizing and reproducing these biases in downstream tasks, with potential societal consequences such as reinforcing political polarization or misrepresenting minority views Feng et al. (2023). While recent research has highlighted the presence of political bias in LLM outputs, the extent to which these biases stem from memorized content in pretraining data remains underexplored.
@@
To address this gap, we propose a pipeline to retrieve relevant documents from the pretraining corpora, then evaluate the political leanings expressed in these documents, and subsequently assess the alignment of political leanings in pretraining corpora with the responses generated by the LLM.
As a case study, we focus on US Supreme Court cases, which frequently address contentious and politically charged issues, such as death penalty, abortion, same-sex marriage, and voting rights, making them strong indicators of political leanings. Leveraging the SCOPE Jessee et al. (2022) [^1] Jessee et al. 2022 has not named the dataset. Hereafter, we refer to the dataset as SCOPE: Supreme COurt Case Political Evaluation. survey data on US Supreme Court cases from political studies, this paper examines the political leanings of eight LLMs and five open-source pre-training corpora, comparing them to human survey responses and Supreme Court rulings. [^2] Our code and data is available at https://github.com/TUMLegalTech/scotus_alignment The main contributions of our work are threefold:
@@
  • Our empirical findings indicate that LLMs exhibit significant alignment with their training corpora, yet we do not find strong alignment with human opinions. This highlights the critical need for methods to detect and mitigate memorized political content in LLMs. We advocate for more transparency in curating training data for LLMs.
@@ 2 Background
2.1 LLMs and their pretraining corpra
@@
Existing studies have explored the impact of biases in training corpora on LLM behavior, primarily through second-stage controlled training setups such as continual pretraining Feng et al. (2023); Chalkidis and Brandl (2024). While continual pretraining can offer valuable insights into the causal links between training data and model outputs, these studies rarely applied to study LLMs’ behavior based on initial pretraining phase, where biases are fundamentally embedded. Additionally, it is also computationally expensive to conduct such extensive continual training experiments on initial phase. An alternative strategy involves investigating the correlation between biases in training corpora and those in model outputs Seshadri et al. (2024). Previously this approach has been underexploited, primarily due to the limited accessibility of large-scale pretraining datasets. Many commercial LLM providers (e.g., GPT-4 by OpenAI 2023 and Claude by Anthropic 2023) disclose minimal information about their training sets, not even corpus size or data source. With the growing call in the academic community for transparency and accessibility of LLM pretraining data Pistilli et al. (2023); McMillan-Major et al. (2024), several organizations have begun to make large-scale pretraining datasets publicly available, including RedPajama (Weber et al., 2024) and Dolma Soldaini et al. (2024). These initiatives are complemented by the development of APIs and analytical tooling platforms, such as WIMDB Elazar et al. (2024), which facilitate comprehensive analysis of the corpora. In this paper, we leverage WIMDB to analyze the political leanings in five publicly accessible corpora and subsequently evaluate how these leanings correlate with the outputs generated by various LLMs.
2.2 Evaluating LLM-Human Alignment
Recent research has increasingly focused on probing LLMs political opinions. Most approaches typically follow a two-stage process: (1) assessing an LLM’s political stance on specific topics, and (2) measuring how closely its responses align with human opinions. A common strategy for evaluating LLM opinions involves using political orientation tests (e.g., Political Compass Test,333www.politicalcompass.org/test as in Röttger et al. 2024; Feng et al. 2023) or survey questionnaires (e.g., PewResearch ATP,444www.pewresearch.org/writing-survey-questions/ as in Santurkar et al. 2023). To quantify the alignment between human and LLM responses, prior work typically measures the similarity of their opinion distributions using either (a) distance-based metrics—such as the 1-Wasserstein distance Santurkar et al. (2023); Sanders et al. (2023) and Jensen–Shannon divergence Durmus et al. (2024)—or (b) statistical analyses, including Cohen’s Kappa Argyle et al. (2023); Hwang et al. (2023) and Pearson correlation coefficients Movva et al. (2024). We refer to Ma et al. 2024 for an extensive survey of methods in this area. In this work, we use SCOPE to probe LLMs’ political opinions because it offers several advantages over the above-mentioned political surveys used in previous studies: The cases in SCOPE are selected by experts, ensuring that they address the most important and publicly salient legal topics. Experts carefully word the questions and response options to be understandable to the general public. Moreover, political experts have annotated each case with thoughtfully chosen keywords, which facilitate our retrieval of relevant documents from large pretraining corpora, as detailed in Sec 4.1.3.
@@ 3.1 Dataset
In political science, researchers often estimate individuals’ or groups’ political preferences and ideological positions by analyzing observable behaviors, such as voting patterns and survey responses Martin and Quinn (2002); Ho et al. (2008). For example, Jessee et al. (2022) created SCOPE to gather respondents’ views on Supreme Court decisions. By comparing collected survey responses with the Court’s voting record, they demonstrated that the Court has adopted a more conservative stance than the general U.S. public. In this study, we use SCOPE to prompt various LLMs to assess their political leanings and subsequently compare their alignment with surveyed human opinions and political leanings in their training data.
The SCOPE dataset comprises 32 cases, each represented by a binary-choice question asking respondents to express their views on the Court’s ruling as either supportive (pro) or opposing (opp). Fig 2 provides an example of a survey question. Tab 2 in App C lists all 32 cases along with their corresponding legal topics in the SCOPE dataset. For each case, between 1,500 and 2,158 respondents indicate whether they are pro or opp regarding the Court’s decision. Additionally, SCOPE captures each respondent’s self-identified political ideology, enabling the categorization of participants into self-identified Democrats or Republicans. Tab 2 in App C showcases the distribution of choices {pro, opp} among the overall surveyed respondents, as well as within the self-identified Democratic and Republican respondents. Further descriptive statistics on respondents’ backgrounds are available in the original study Jessee et al. (2022).
@@
Table 1: Overview of evaluated LLMs, along with their pretraining dataset. * signifies that the model was not trained exactly on this dataset, due to filtering, using additional data, or the original data being private.
@@ 3.3 Pretraining Corpora
Tab 1 lists the corresponding pretraining corpora (when available) of the LLMs we investigated in this work. It is important to note that among the various pairs of LLM and their pretraining corpora we consider, only the OLMo-SFT and OLMo-Instruct models were trained directly on the pretraining corpus Dolma Soldaini et al. (2024). While for all other pairs, the LLMs may not have been trained exactly on the versions of the corpora we consider, due to factors such as filtering, or inclusion of additional data Elazar et al. (2024). Despite these discrepancies, we treat the documented corpora as reasonable proxies for analysis, as they represent the closest publicly available approximations of the actual training data for these models. [^5] All models examined in this paper have undergone post-training, such as Supervised or Instruction Fine-Tuning, which may also influence the opinions in models outputs. However, prior research Feng et al. (2023) suggests that the shift introduced by post-training is relatively small. We also explored the correlation between LLMs’ political leanings and that in their post-training data, but did not observe any significant correlation. Further discussions can be found in App G.
@@ 4 Methodology
We employ a three-stage process to examine LLMs’ political leanings and compare their alignment with surveyed human opinions and their pretraining corpora, whenever available. First, we introduce how we assess the political leanings of different entities. [^6] We use entity to refer to either a group of surveyed respondents, Supreme Court justices, LLM-generated responses, or content within the training data. Next, we measure the political leanings alignment among them in Sec 4.2. Finally, we conduct significance tests to determine whether the observed differences in LLM alignment with different entities are statistically significant in Sec 4.3.
@@ 4.1 Extracting the Preference Distributions
In this section, we outline the methodology used to extract the preference distribution of various entities. We divide the entities to three categories - humans $D_{H}$, LLMs $D_{M}$, and pretraining corpora $D_{C}$.
Figure 4: For each survey case in SCOPE, we created six different prompt templates, and we then sample five responses from each of the six prompt variations, yielding in a total of 30 responses per case per model.
4.1.1 Humans
Under the category of humans, we consider the preference distribution of four entities: $D_{H}=\{D_{\mathrm{pub}},D_{\mathrm{dem}},D_{\mathrm{rep}},D_{\mathrm{court}}\}$. Here, $D_{\mathrm{pub}}$ represents the preference distribution of the overall surveyed respondents, while $D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$ correspond to surveyed self-identified Democrat and Republican respondents, respectively; $D_{\mathrm{court}}$ represents the preference distribution of the Court. All preference distributions are Bernoulli, with the respective parameter estimated from the data. For the $D_{\mathrm{court}}$, we fetch the judges’ votes from the Supreme Court Database Spaeth et al. (2024), [^7] http://scdb.wustl.edu/ and then calculate $D_{\mathrm{court}}$ as the ratio of justices who agree (pro) / dissent (opp) with the majority decision. For $D_{pub},D_{\mathrm{dem}}$ and $D_{\mathrm{rep}}$ we calculate them as the ratios of {pro} versus {opp} to the court’s decisions among the respondents based on data retrieved from SCOPE.
4.1.2 LLMs
Under the category of the LLMs $D_{M}$, we probe the political preferences of eight LLMs as listed in Tab 1. Following Scherrer et al. 2023, for each survey case in SCOPE, we created six different prompt templates, as illustrated in Fig 4. We then sample five responses from each of the six prompt variations from the LLMs at a temperature setting of 1, yielding in a total of 30 responses per case per model. The complete prompt templates and detailed prompt creation process can be found in 7(c) in App B.
To map the LLM-generated answers to one of the given choice options, we employed an iterative, rule-based matching pipeline, as explained in App B, as illustrated in Fig 4. The preference distribution, denoted as $D_{m}=p_{m}(a_{j}\mid q_{i})$, reflects the ratio of support versus opposition to the court’s decision across the 30 generated responses for model $m$ on case $q_{i}$.
4.1.3 Pretraining Data
@@
For each of the 32 cases $q_{i}$ in the SCOPE survey, we compile a set of keywords $K_{i}$ to retrieve relevant documents $T_{i}$ from the pretraining corpora using the WIMDB API Elazar et al. (2024), a tool designed to facilitate analysis of large-scale pretraining corpora. For example, in the case Baze v. Rees, [^8] Baze v. Rees, 553 U.S. 35 (2008), addresses whether lethal injection for executions was constitutional or not. we use keywords such as [lethal injection; capital punishment; Baze; Rees] retrieving 206 documents from the Dolma corpus. Further details on keyword selection and retrieval statistics for each case are provided in App C. Additionally, an example of a retrieved document is included in App I.
@@
We use zero-shot Llama3-70B to assess the political stance $s_{i}^{j}$ of each retrieved document $t_{i}^{j}\in T_{i}$. The model is prompted to evaluate the document’s level of support for the court’s decision on a Likert scale from 1 (strongly against) to 5 (strongly supportive). If a document is unrelated to the case’s political issue, the model is instructed to return ‘Not related’. The complete prompt template we sed to evaluate the stance scores of the retrieved documents is available in Fig 8 in App B.
@@
Quality Assessment of Stance Detection To evaluate the reliability of Llama3-70B’s stance detection, we manually annotated a randomly selected sample of 80 retrieved documents. We measure the agreement between human and model labels using Spearman’s rank correlation Spearman (1904). The overall Spearman’s $\rho$ is 0.68, indicating good alignment between Llama370B and human annotators. App C offers details on the quality assessment process. To evaluate the robustness of our document retrieval method, we performed a bootstrapping analysis by iteratively excluding 20% of retrieved documents. This procedure revealed no significant shifts in measured political leanings (see App C for methodological details). Although differences in keyword selection may affect document retrieval and thereby influence corpus-level political stance estimates, our findings demonstrate that results are resilient to changes in the retrieved documents.